OpenStudy (anonymous):

(x^3-y^2)^13 find the x^18th term

OpenStudy (whpalmer4):

Definitely a problem where working smarter is better!

OpenStudy (whpalmer4):

Use the binomial theorem: \[(a+b)^n = a^n + na^{n-1}b + \frac{n(n-1)}{2}a^{n-2}b^2 + ... + b^n\]

OpenStudy (anonymous):

Alright let me give this a try

OpenStudy (whpalmer4):

What did you get?